OpenStudy (anonymous):

Describe the difference between an angle with a positive and an angle with a negative measure

OpenStudy (anonymous):

This is trig?

OpenStudy (anonymous):

^you gonna delete that now, lagrange? trololo.

OpenStudy (anonymous):

wow man chill out, whats your deal

OpenStudy (anonymous):

precaculus

OpenStudy (anonymous):

okay, well then we can talk about positive and negative angles with reference to the unit circle

OpenStudy (anonymous):

my deal is you making smartass comments on every question, which wastes everyones time. How about getting straight to the answer?

OpenStudy (anonymous):

kthanksbye.

OpenStudy (anonymous):

anyway, we can think of an angle basically as a rotation of a line that begin on the x axis

OpenStudy (anonymous):

now, if we think of that this way, we logically conclude that a positive angle increases through the 1st, 2nd, 3rd and 4th quadrents

OpenStudy (anonymous):

now what can we conclude of about a negative angle, well a negative angle would in turn decrease throught the 1st, 2nd , 3rd and 4th quadrents

OpenStudy (anonymous):

i mean, decrease through the 4th, 3rd, 2nd and 1st quadrents

OpenStudy (anonymous):

is that okay?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

The negative simply tells you that you're rotating the angle clockwise (opposite to normal). \[-\theta = 360 - \theta\]

OpenStudy (anonymous):

Really need pictures to illistrate :S

OpenStudy (anonymous):

and positive tells you that you are rotating counterclockwise

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

This picture illistrates the differences...

OpenStudy (anonymous):

illustrate even :)

OpenStudy (anonymous):

what I say? + , _ angle?

OpenStudy (anonymous):

My answer would be what i said in my first reply. Well i'd mention the cartesian plane too. How i'd answer it... "The negative sign tells you that the angle is being rotated clockwise (opposite to normal) through the cartesian plane. \[-\theta = 360 - \theta\]"